OpenStudy (anonymous):

I give medals! Please help! A hectare of land yields roughly 370 bushels of corn (150 bushels per acre). A bushel consists of 1250 ears of corn, and each ear contains 80 food calories of energy. Assume a person eats only corn and requires 2000 calories per day. Ignore the corn thing and pretend its now beef. It takes about 20 times as much land (compared to corn) to raise beef calories (land is needed to raise the corn to feed the cows.) How much land would be needed to feed the person if they ate only beef?

OpenStudy (btaylor):

This is a good problem. First, we need to find the amount of land that is required to feed 1 person: \[\frac{2000 calories}{} \times \frac{1 ear}{80 calories} \times \frac{1 bushel}{1250 ears} \times \frac{1 hectare}{370 bushels}=0.000054054 hectares\] \[0.000054054 hectares \times 20 = 0.00108108108 hectares\]